# Robust AC vector sensing at zero magnetic field with pentacene

### v1

@@ Abstract
Quantum sensors based on electronic spins have emerged as powerful probes of microwave-frequency fields. Among other solid-state platforms, spins in molecular crystals offer a range of advantages, from high spin density to functionalization via chemical tunability. Here, we demonstrate microwave vector magnetometry using the photoexcited spin triplet of pentacene molecules, operating at zero external magnetic field and room temperature. We achieve full three-dimensional microwave field reconstruction by detecting the Rabi frequencies of anisotropic spin-triplet transitions associated with two crystallographic orientations of pentacene in deuterated naphthalene crystals. We further introduce a phase alternated protocol that extends the rotating-frame coherence time by an order of magnitude and enables sensitivities of approximately $1~\mu\mathrm{T}/\sqrt{\mathrm{Hz}}$ with sub-micrometer spatial resolution. These results establish pentacene-based molecular spins as a practical and high-performance platform for microwave quantum sensing in addition to demonstrating control techniques broadly applicable to other molecular and solid-state spin systems.
@@ I Introduction
Quantum sensors based on optically-detected solid-state defect spin qubits have demonstrated remarkable sensitivity to electromagnetic fields Taylor et al. (2008); Maze et al. (2008); Pham et al. (2011); Neumann et al. (2013); Wang et al. (2023), enabling nanoscale imaging and precision metrology for spin-wave imaging Lee-Wong et al. (2020); Van der Sar et al. (2015), micro-electronic device characterization Lillie et al. (2019); Nowodzinski et al. (2015), and bio-sensing Tan et al. (2022); Miller et al. (2020). Despite these advances, defect-based quantum sensors face intrinsic limitations. Their scalability is constrained by the difficulty of creating high-density, uniform defect ensembles, as the accompanying increase in surrounding spin impurities leads to decoherence Zhou et al. (2020); Ghassemizadeh et al. (2024); Bauch et al. (2020); Park et al. (2022). Furthermore, at low external magnetic fields, inhomogeneous charge and strain environments mix the spin energy levels, degrading both the coherence and the sensing accuracy Mittiga et al. (2018).
Molecular quantum systems Köhler et al. (1993); Wrachtrup et al. (1993) with optically addressable spin degrees of freedom have recently emerged as a compelling alternative route for quantum sensing Singh et al. (2025a); Mena et al. (2024); Singh et al. (2025b); Yamauchi et al. (2024). These molecular spins can exhibit long coherence times and combine the advantages of chemically programmable spin sites with high density doping in many host crystals Zadrozny et al. (2017); Bayliss et al. (2020). These properties allow molecular color centers to operate under conditions where conventional solid-state defect qubits typically fail, such as in flexible substrates or amorphous environments Pei (2022). Optical addressability further enables efficient initialization and read-out of spin states.
@@
Among these systems, pentacene stands out for its long-lived photoexcited triplet states and efficient optical spin polarization at room temperature Singh et al. (2025a); Mena et al. (2024); Köhler et al. (1993); Wrachtrup et al. (1993). As an organic molecule, pentacene can be incorporated into thin films Moro et al. (2022); Lin et al. (1997), patterned directly on chip Yunus et al. (2022), or grown as bulk crystals, offering versatility for device integration and scalable sensing architectures.
@@
Here, leveraging an ensemble of photoexcited triplet states of pentacene molecules embedded in a naphthalene single crystal, we develop a Rabi-based protocol that exploits the intrinsic multi-orientational structure and anisotropic spin transitions for vector AC magnetic-field sensing at microwave frequencies. To further enhance its sensitivity, we introduce a phase-modulated control scheme Solomon (1959); Aiello et al. (2013); Hirose et al. (2012) that effectively decouples the qubit from field fluctuations and inhomogeneities that arise from the environment and the driving itself. With this approach, we demonstrate magnetic-field sensitivities reaching $1~\mu\mathrm{T}/\sqrt{\mathrm{Hz}}$ under ambient conditions with sub-micrometer spatial resolution. Overall, these results establish a viable pathway toward room-temperature quantum sensing and vector-field imaging using molecular spin systems.
@@ II.1 Pentacene sample and spin properties
We used high-quality pentacene-doped naphthalene single crystals with a pentacene concentration of about $4\times 10^{-5}$ mol/mol Quan et al. (2023) (see supplementary materials sup for detailed sample preparation and characterization). In pentacene doped into naphthalene crystals, the pentacene molecules align along two crystallographically distinct orientations from substitution of the naphthalene host at inequivalent lattice sites, as shown in Fig. 1(a). These orientations are fixed relative to the crystal axes, thereby establishing a well-defined geometric relation between the molecular frames and the laboratory sensing frame [Fig. 1(b)].
@@
At zero external magnetic field, the two molecular orientations of pentacene are energetically degenerate. The simplified energy-level structure of an individual pentacene molecule in Fig. 1(d) Singh et al. (2025a); Mena et al. (2024) highlights the presence of a spin-triplet metastable state. Under optical excitation, population can be pumped from the singlet ground state ($S_{0}$) to the excited singlet state ($S_{1}$), from which it can either decay radiatively back to $S_{0}$ or undergo spin-dependent intersystem crossing (ISC) into the triplet manifold ($T_{X}$, $T_{Y}$, $T_{Z}$). The spin-dependent ISC process preferentially populates the $T_{X}$ state, enabling optical spin polarization Wu et al. (2019). To prepare an initial triplet polarized on other sublevels, a fast ($\sim$100 ns) flip microwave pulse is employed immediately after laser pumping.
Because the decay rates $\Gamma_{\mu}$ ($\mu=X,Y,Z$) from the triplet sublevels to the ground state $S_{0}$ are unequal, the population of the $S_{0}$ state after a delay time $t_{d}$ encodes information about the previous excited spin state. The resulting photoluminescence (PL), which is proportional to the population in $S_{0}$, thus provides spin-dependent contrast and enables optical readout of the triplet manifold.
@@
with zero-field splitting parameters $D=1389~(2\pi)\mathrm{MHz}$ and $E=-48~(2\pi)\mathrm{MHz}$, which need to be experimentally determined for each host matrix. Transitions between the triplet eigenstates can be driven by a microwave field $\vec{B}_{\mathrm{ac}}(t)\!\cdot\!\vec{S}$ through the nonzero matrix elements
@@
where the stretch exponents $\iota_{\mu,\nu}$ capture the multichannel relaxation pathways. This defines the readout contrast $C(t_{d})$. Experimentally, $t_{d}=30~\mu\mathrm{s}$ and $47~\mu\mathrm{s}$ maximize the contrast for the $T_{X}\leftrightarrow T_{Z}$ and $T_{Y}\leftrightarrow T_{Z}$ qubits, respectively.
Figure 1(e) shows the optically detected magnetic resonance (ODMR) of the $T_{X}\leftrightarrow T_{Z}$ ($\omega_{xz}=D-E$) and $T_{Y}\leftrightarrow T_{Z}$ ($\omega_{yz}=D+E$) transitions at zero magnetic field.
@@ II.2 Principle and experimental demonstration
A more compact protocol can be obtained exploiting the different pentacene orientations in the crystal, similar to prior schemes with NVs Wang et al. (2015); Schloss et al. (2018). As discussed above, the two pentacene orientations share a common molecular $X$-axis but have distinct $Y$-axes ($Y_{1}$ and $Y_{2}$). Therefore, when an AC microwave field is applied, the $T_{Y}\leftrightarrow T_{Z}$ transition in both orientations is driven equally by the $X$-component of the field, whereas the $T_{X}\leftrightarrow T_{Z}$ transition probes the orientation-dependent $y$-components ($Y_{1}$ and $Y_{2}$), as illustrated in the inset of Fig. 2(d), similarly to prior schemes with NVs Schloss et al. (2018); Clevenson et al. (2018). Consequently, Rabi measurement of $T_{X}\leftrightarrow T_{Z}$ transitions yields two Rabi frequencies ($\Omega_{Y_{1}}$, $\Omega_{Y_{2}}$) from the two orientations [Fig. 2(b,d)] and one ($\Omega_{X}$) from the $T_{Y}\leftrightarrow T_{Z}$ transition [Fig. 2(c,d)]. This three-axis projection scheme enables full reconstruction of the three-dimensional AC magnetic-field vector. For simplicity, we define the orthogonal laboratory sensing frame $XY^{\prime}Z^{\prime}$, where the $Y^{\prime}$ axis bisects the molecular $Y_{1}$ and $Y_{2}$ axes [Fig. 1(b) and Fig. 2(d), inset] and $Z^{\prime}$-axis aligns with the $b$ axis of naphthalene lattice. The AC magnetic-field components in this frame can be reconstructed as
@@ III.1 Decoherence sources in driven pentacene spin systems
When the goal is to sense weak AC fields, $\xi_{\Omega_{1}}$ can be neglected and longitudinal inhomogeneous broadening limits the coherence time. In deuterated pentacene, the observed $\sim 0.6~(2\pi)\mathrm{MHz}$  linewidth is dominated by hyperfine coupling to the deuterons in the pentacene molecule Singh et al. (2025a); Quan (2021). This coupling introduces a distribution of quasi-static detunings $\delta\omega$ that yield an effective Rabi frequency $\Omega_{R}(\delta\omega)=\sqrt{\Omega_{1}^{2}+\delta\omega^{2}}.$ At low microwave amplitudes ($\Omega_{1}\lesssim\delta\omega$), the finite linewidth introduces a bias in the Rabi oscillation frequency, faster dephasing, and reduces the contrast sup , thereby degrading the sensing accuracy Billaud et al. (2025); Kitamura et al. (2025). A simple strategy to mitigate the broadening effects would be to add a bias driving field to the target AC, in order to operate at larger driving strengths. When $\langle\Omega_{R}\rangle\approx\Omega_{1}+{\langle\delta\omega^{2}\rangle}/{2\Omega_{1}}$, detuning-induced decoherence and bias is strongly suppressed.
@@ IV Discussion and Outlook
Experimental evidence indicates an inherent bandwidth–sensitivity trade-off for standard Rabi sensing. The longitudinal noise limits the sensitivity in low field regime, while the inhomogeneity of the strong drive introduces additional dephasing. The rotary echo sequence applies strong driving with periodic phase inversions which refocuses the inhomogeneous drive amplitudes. This provides a perfect biasing of the weak field to the strong driving regime, thus effectively decouples to the noise field. The resulting lifetime of the dressed-state coherence is significantly extended in an order of magnitude, enabling accurate, broadband estimation of weak AC fields. The coherence enhancement provided by RE is particularly important for pentacene-based protocols, which necessitates a substantial overhead time to build spin-state contrast, causing the overall sensitivity to scale linearly with the available coherence.
@@
An important advantage of using pentacene molecular spins to perform AC magnetometry is the possibility to work at zero external magnetic field, as we demonstrated. In this regime, the accuracy and sensitivity of traditional solid-state defect sensors degrade due to local strain, charge and external defects that induce state mixing and decoherence Vetter et al. (2022); Zheng et al. (2019); Udvarhelyi et al. (2018); Ungar et al. (2025); Mittiga et al. (2018). Additionally, the triplet level structure of pentacene is intrinsically tolerant to small background magnetic fields, whose influence is quadratically suppressed, (e.g., producing less than $20~(2\pi)\,\mathrm{kHz}$ resonance shifts and negligible state mixing under Earth magnetic field). This insensitivity to weak fields makes pentacene a robust and practically convenient platform in unshielded or imperfectly controlled environments.
To achieve better performance in quantum sensing, molecular spin systems offer unique opportunities for chemical engineering, including nuclear-spin Atzrodt et al. (2018); Ryan et al. (2025) or ligand modification Graham et al. (2017); Mirzoyan et al. (2021); Wedge et al. (2012), as well as programmable spatial placement of the spin hosts Lavroff et al. (2021); Li et al. (2025). The former allows tailoring the spin properties; for example, further deuteration of the naphthalene host can suppress nuclear-spin noise and extend coherence Salvadori et al. (2017); Bayliss et al. (2022); Li et al. (2025). The latter enables precise positioning of molecular spins, for instance through metal–organic framework architectures Falcaro et al. (2014); Yamabayashi et al. (2018); Yamauchi et al. (2022), opening pathways toward controllable nonlocal correlation sensing.
These advantages have long motivated interest in exploring molecular spin–qubit platforms, pursuing better coherence properties, enhanced robustness and more environmentally suitable form factors. Our work bridges these principles into practice, establishing a powerful and scalable approach for room-temperature quantum sensing.
@@ V Acknowledgment
We thank Ashok Ajoy, Shimon Kolkowitz, Andrew Stasiuk, Bo Xing, Keyuan Zhong for helpful discussions. This work was supported in part by the National Science Foundation under Grants No. PHY2317134 (Center for Ultracold Atoms). B. L. thanks MathWorks for their support in the form of a Graduate Student Fellowship. The opinions and views expressed in this publication are from the authors and not necessarily from MathWorks.

### v2

@@ Abstract
Quantum sensors based on electronic spins have emerged as powerful probes of microwave-frequency fields. Among other solid-state platforms, spins in molecular crystals offer a range of advantages, from high spin density to functionalization via chemical tunability. Here, we demonstrate microwave vector magnetometry using the photoexcited spin triplet of deuterated pentacene molecules, operating at zero external magnetic field and room temperature. We achieve full three-dimensional microwave field reconstruction by detecting the Rabi frequencies of anisotropic spin-triplet transitions associated with two crystallographic orientations of pentacene in naphthalene crystals. We further introduce a phase alternated protocol that extends the rotating-frame coherence time by an order of magnitude and enables sensitivities of $1~\mu\mathrm{T}/\sqrt{\mathrm{Hz}}$ with sub-micrometer spatial resolution. These results establish pentacene-based molecular spins as a practical and high-performance platform for microwave quantum sensing, and the control techniques are broadly applicable to other molecular and solid-state spin systems.
@@ I Introduction
Quantum sensors based on optically-detected solid-state defect spin qubits have demonstrated remarkable sensitivity to electromagnetic fields Taylor et al. (2008); Maze et al. (2008); Pham et al. (2011); Neumann et al. (2013); Wang et al. (2023), enabling nanoscale imaging and precision metrology for spin-wave imaging Lee-Wong et al. (2020); Van der Sar et al. (2015), micro-electronic device characterization Lillie et al. (2019); Nowodzinski et al. (2015), and bio-sensing Tan et al. (2022); Miller et al. (2020). Despite these advances, defect-based quantum sensors face intrinsic limitations. Their scalability is constrained by the difficulty of creating high-density, uniform defect ensembles, as the accompanying increase in surrounding spin impurities leads to decoherence Zhou et al. (2020); Ghassemizadeh et al. (2024); Bauch et al. (2020); Park et al. (2022); Mittiga et al. (2018); Wang et al. (2022). Furthermore, at low external magnetic fields, inhomogeneous charge and strain environments mix the spin energy levels, degrading both the coherence and the sensing accuracy Zhu et al. (2014); Matsuzaki et al. (2016, 2015).
@@
Molecular quantum systems Köhler et al. (1993); Wrachtrup et al. (1993) with optically addressable spin degrees of freedom have recently emerged as a compelling alternative route for quantum sensing Singh et al. (2025); Mena et al. (2024); Singh et al. (2025); Yamauchi et al. (2024). These molecular spins can exhibit long coherence times and combine the advantages of chemically programmable spin sites with high density doping in many host crystals Zadrozny et al. (2017); Bayliss et al. (2020). Among these systems, pentacene stands out for its long-lived photoexcited triplet states and efficient optical spin polarization at room temperature Singh et al. (2025); Mena et al. (2024); Köhler et al. (1993); Wrachtrup et al. (1993). Recent studies incorporate pentacene into thin films Moro et al. (2022); Lin et al. (1997), patterned directly on chip Yunus et al. (2022), or grown as bulk crystals, demonstrating versatility for device integration and scalable sensing architectures Zhang et al. (2024); Wu et al. (2022).
@@
Here, leveraging an ensemble of photoexcited triplet states of pentacene molecules embedded in a naphthalene single crystal, we develop a Rabi-based protocol that exploits the intrinsic multi-orientational structure and anisotropic spin transitions for vector AC magnetic-field sensing at microwave frequencies. To further enhance its sensitivity, we introduce a phase-modulated control scheme Solomon (1959); Aiello et al. (2013); Hirose et al. (2012) that effectively decouples the qubit from field fluctuations and inhomogeneities that arise from the environment and the driving itself. The magnetic-field sensitivities reaches $1~\mu\mathrm{T}/\sqrt{\mathrm{Hz}}$ with sub-micrometer spatial resolution.
@@ II.1 Pentacene sample and spin properties
We used pentacene-d14-doped naphthalene single crystals with a pentacene concentration of about $4\times 10^{-5}$ mol/mol Quan et al. (2023). The sample was extensively purified to achieve high quality single crystals and minimize impurities (see supplementary materials sup for detailed sample preparation and characterization). The pentacene molecules align along two crystallographically distinct orientations from substitution of the naphthalene host at inequivalent lattice sites [Fig. 1(a)]. These orientations are fixed relative to the crystal axes, thereby establishing a well-defined geometric relation between the molecular frames and the laboratory sensing frame [Fig. 1(b)].
@@
At zero external magnetic field, the two oriented of pentacene molecules are energetically degenerate. An electronic spin-triplet metastable state Singh et al. (2025); Mena et al. (2024) presents in a pentacene molecule, as shown in the simplified energy-level structure [Fig. 1(d)]. Under optical excitation, population can be pumped from the singlet ground state ($S_{0}$) to the excited singlet state ($S_{1}$), from which it can either decay radiatively back to $S_{0}$ or undergo spin-dependent intersystem crossing (ISC) into the triplet manifold ($T_{X}$, $T_{Y}$, $T_{Z}$). The spin-dependent ISC process preferentially populates the $T_{X}$ state, enabling optical spin polarization. To prepare an initial triplet polarized on other sublevels, a fast ($\sim$100 ns) flip microwave pulse is employed immediately after laser pumping.
The triplet state decay to the ground state $S_{0}$ also via spin-dependent ISC, with decay rates $\Gamma_{\mu}$ ($\mu=X,Y,Z$) that are unequal among the triplet sublevels. Therefore, the population of the $S_{0}$ state after a delay time $t_{d}$ encodes information about the previous excited spin state. The resulting photoluminescence (PL), which is proportional to the population in $S_{0}$, thus provides spin-dependent contrast and enables optical readout of the triplet manifold.
@@
with zero-field splitting parameters $D=1389~(2\pi)\mathrm{MHz}$ and $E=-48~(2\pi)\mathrm{MHz}$ Quan (2021). Transitions between the triplet eigenstates can be driven by a microwave field $\vec{B}_{\mathrm{ac}}(t)\!\cdot\!\vec{S}$ through the nonzero matrix elements
@@
where the stretch exponents $\iota_{\mu,\nu}$ capture the multichannel relaxation pathways. This defines the readout contrast $C(t_{d})$. Experimentally, $t_{d}=30~\mu\mathrm{s}$ and $47~\mu\mathrm{s}$ maximize the contrast for the $T_{X}\leftrightarrow T_{Z}$ and $T_{Y}\leftrightarrow T_{Z}$ qubits, respectively. Figure 1(e) shows the optically detected magnetic resonance (ODMR) of the two transitions at zero magnetic field. More details are presented in Supplementary materials.
[…]
@@ II.2 Principle and experimental demonstration
A more compact protocol can be obtained exploiting the different pentacene orientations in the crystal, similar to prior schemes with NVs Wang et al. (2015); Schloss et al. (2018). As mentioned above, the two pentacene orientations share a common molecular $X$-axis but have distinct $Y$-axes ($Y_{1}$ and $Y_{2}$). Therefore, when an AC microwave field is applied, the $T_{Y}\leftrightarrow T_{Z}$ transition in both orientations is driven equally by the $X$-component of the field, whereas the $T_{X}\leftrightarrow T_{Z}$ transition probes the orientation-dependent $y$-components ($Y_{1}$ and $Y_{2}$), as illustrated in the inset of Fig. 2(d). Consequently, Rabi measurement yields two Rabi frequencies ($\Omega_{Y_{1}}$, $\Omega_{Y_{2}}$) from $T_{X}\leftrightarrow T_{Z}$ transitions in two orientations [Fig. 2(b,d)], and one ($\Omega_{X}$) from the $T_{Y}\leftrightarrow T_{Z}$ transition [Fig. 2(c,d)]. This three-axis projection scheme enables full reconstruction of the three-dimensional AC magnetic-field vector. For simplicity, we define the orthogonal laboratory sensing frame $XY^{\prime}Z^{\prime}$, where the $Y^{\prime}$ axis bisects the molecular $Y_{1}$ and $Y_{2}$ axes [Fig. 1(b) and Fig. 2(d), inset] and $Z^{\prime}$-axis aligns with the $b$ axis of naphthalene lattice. The AC magnetic-field components in this frame can be reconstructed as
@@ III.1 Decoherence sources in driven pentacene spin systems
When the goal is to sense weak AC fields, $\xi_{\Omega_{1}}$ can be neglected and longitudinal inhomogeneous broadening limits the coherence time. In deuterated pentacene, the observed $\sim 0.6~(2\pi)\mathrm{MHz}$  linewidth is dominated by hyperfine coupling to the deuterons in the pentacene molecule Singh et al. (2025); Quan (2021). This coupling introduces a distribution of quasi-static detunings $\delta\omega$ that yield an effective Rabi frequency $\Omega_{R}(\delta\omega)=\sqrt{\Omega_{1}^{2}+\delta\omega^{2}}.$ At low microwave amplitudes ($\Omega_{1}\lesssim\delta\omega$), the finite linewidth introduces a bias in the Rabi oscillation frequency, faster dephasing, and reduces the contrast sup , thereby degrading the sensing accuracy Billaud et al. (2025); Kitamura et al. (2025). A simple strategy to mitigate the broadening effects would be to add a bias driving field to the target AC, in order to operate at larger driving strengths. When $\langle\Omega_{R}\rangle\approx\Omega_{1}+{\langle\delta\omega^{2}\rangle}/{2\Omega_{1}}$, detuning-induced decoherence and bias is strongly suppressed.
@@ IV Discussion and Outlook
Experimental evidence indicates an inherent bandwidth–sensitivity trade-off for standard Rabi sensing. The longitudinal noise limits the sensitivity in low field regime, while the inhomogeneity of the strong drive introduces additional dephasing. The RE sequence applies strong driving with periodic phase inversions which refocuses the inhomogeneous drive amplitudes, thus providing a perfect biasing of the weak field to the strong driving regime and effectively decouples to the noise field. The coherence enhancement is particularly important for pentacene-based protocols, which necessitates a substantial overhead time to build spin-state contrast, causing the overall sensitivity to scale linearly with the available coherence. Recent advances in rotating-frame coherence protection for NV centers  Stark et al. (2017); Salhov et al. (2024); Kitamura et al. (2025) may inspire future studies on their applicability to pentacene.
@@
An important advantage of using pentacene molecular spins to perform AC magnetometry is the possibility to work at zero external magnetic field. In this regime, the accuracy and sensitivity of typical solid-state defect sensors degrade due to local strain, charge and external defects that induce state mixing and decoherence Vetter et al. (2022); Zheng et al. (2019); Udvarhelyi et al. (2018); Ungar et al. (2025); Mittiga et al. (2018); Zhu et al. (2014); Matsuzaki et al. (2016, 2015); Saijo et al. (2018). In contrast, the vast majority of impurities and lattice defects can be removed through extensive purification of organic crystals. Additionally, the triplet level structure of pentacene is intrinsically robust to small background magnetic fields, whose influence is quadratically suppressed (e.g., $20~(2\pi)\,\mathrm{kHz}$ resonance shifts under Earth magnetic field). This insensitivity makes pentacene a practically convenient platform in unshielded or imperfectly controlled environments.
Looking forward, molecular spin systems offer unique opportunities for quantum applications, benefiting from chemical engineering, including nuclear-spin Atzrodt et al. (2018); Ryan et al. (2025) or ligand modification Graham et al. (2017); Mirzoyan et al. (2021); Wedge et al. (2012), as well as programmable spatial placement of the spin hosts Lavroff et al. (2021); Falcaro et al. (2014); Yamabayashi et al. (2018); Yamauchi et al. (2022). These advantages have long motivated interest in exploring molecular spin–qubit platforms, pursuing better coherence properties, enhanced robustness and more environmentally suitable form factors. Our work bridges these principles into practice, establishing a powerful and scalable approach for room-temperature quantum sensing.
@@ V Acknowledgment
We thank Ashok Ajoy, Shimon Kolkowitz, Max Attwood, Andrew Stasiuk, Bo Xing, Keyuan Zhong and Takuya Isogawa for helpful discussions. This work was supported in part by the National Science Foundation under Grants No. PHY2317134 (Center for Ultracold Atoms). B. L. thanks MathWorks for their support in the form of a Graduate Student Fellowship. The opinions and views expressed in this publication are from the authors and not necessarily from MathWorks.